The president of a company was interested in determining whether there is a correlation between sales made by different sales teams and hours spent on employee training. These figures are shown.

Sales
(in thousands)

Training
Hours

11 7
33 11
27 10
38 15
8 3
  1. Compute the correlation coefficient for the data. (Round your answer to 4 decimal places, the tolerance is +/-0.0001.)
    The correlation coefficient is

2. Using the data, what would you expect sales to be if training was increased to eighteen hours? Use the linear regression model. (Round your answer to 2 decimal places, the tolerance is +/-0.01.)
Sales = . (in thousands)
Answer #1

Let the training hours be denoted by x and sales by y

Let the Regression line be y = bo + b1x

where,

bo = ( Σy Σx2 - Σx Σxy ) / ( nΣx2 - (Σx)2 )

b1 = ( nΣxy - ΣxΣy ) / ( nΣx2 - (Σx)2 )

Sales(y) Training Hours(x) x2 xy y2
11 7 49 77 121
33 11 121 363 1089
27 10 100 270 729
38 15 225 570 1444
8 3 9 24 64
Σ 117 46 504 1304 3447

=> bo = ( 117*504 - 46*1304 ) / ( 5*504 - 462 ) = -2.51

b1 = ( 5*1304 - 46*117 ) / ( 5*504 - 462 ) = 2.82

=> y = -2.51 + 2.82x

Correlation coefficient r = ( nΣxy - ΣxΣy ) / sqrt(( nΣx2 - (Σx)2 ) ( nΣy2 - (Σy)2 ))

=> r = ( 5*1304 - 46*117 ) / sqrt(( 5*504 - 462 ) ( 5*3447 - 1172 )) = 0.9508

2. We found the regression line y = -2.51 + 2.82x

when training hours x = 18,

Sales y = -2.51 + 2.82*8 = 20.05